OpenStudy (kayc):

When f(x) is divided by ( x-3 ), the remainder is 2. So if i divided 2[f(x)]^2-3 by ( x-3 ) what is the remainder?

OpenStudy (raden):

because f(3)=2, then f(3)+5=2+5=7 i sure, u understand what i mean :)

OpenStudy (kayc):

haha, yes thank you

OpenStudy (raden):

welcome :)

OpenStudy (fellowroot):

@RadEn I really like your idea, but I think you are missing the part where you are dividing one into another. But its true that if you take the remainder and then add 5 to it you get 7. Works the same with mine.

OpenStudy (raden):

ok, thank u for remember it